Solve the equation
-{6x+(4x+3)]=7-(5x+6)

Answer by amarjeeth123(569)

-{6x+(4x+3)]=7-(5x+6)
-(10x+3)=-5x+1
10x+3=5x-1
5x=-4
x=-4/5=-0.8